OpenStudy (anonymous):

x + y = 5 x = 4y please show the steps

Parth (parthkohli):

Refer to this post by YOU. http://openstudy.com/study#/updates/4fe1edd5e4b06e92b870d45c

OpenStudy (jackellyn):

If you substitute 4y into the first equation and simplify you can solve for y Once you solve for y, substitute into second equation for x (4y) + y =5 5y = 5 y =? x=4(?) x=??

OpenStudy (jackellyn):

Understand?

OpenStudy (anonymous):

i am not seeing it

OpenStudy (jackellyn):

Let's work on the first equation They give you x =4y x +y = 5 (4y) + y =5 because we substitute 4y for x Then add 4y and y 5y = 5 Divide both sides by 5 to get the y alone 5y/5 = 5/5 what is 5/5?

Parth (parthkohli):

Erm......?

OpenStudy (anonymous):

its one

OpenStudy (jackellyn):

so y =1 Then you can put that in the second equation x=4(1) 4*1 is ?

OpenStudy (anonymous):

is 4

OpenStudy (jackellyn):

So x =4 and y=1 To check substitute back into any equation 4+1 = 5 Is that right?

OpenStudy (anonymous):

right

OpenStudy (jackellyn):

That's all then.

OpenStudy (anonymous):

sweet ok i will use that to get the answer for my actual problem i try to put in things simalar to the actual problem so i can figure it out

OpenStudy (jackellyn):

good :)